OpenStudy (anonymous):

Do the ratios 8/9 and 16/18 form a proportion?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Double the first one..

OpenStudy (anonymous):

yes, the do as 8*18=16*9=144